My sister is 15 years older than I am. The sum of our age is 49. Find our age
Answer by ikleyn(52776)

.
Let x = "my" age.

Then the sister's age = x+15.


The sum of ages equation is


    x + (x+15) = 49.


    2x + 15 = 49

    2x      = 49 - 15 = 34.

     x                = 34/2 = 17.


ANSWER. My age is 17.  The sister's age is 17+15 = 32.


CHECK.  17 + 32 = 49.   ! Correct !

Solved.